OpenStudy (anonymous):

ln(-5x+3)= ln 2x+2. I have -5x+3/2x=e^2 to start, am i on the right track?

OpenStudy (tkhunny):

One must first establish what it is that you mean. Is it this \(ln(-5x+3) = ln(2x+2)\)?

OpenStudy (anonymous):

Hi friend, since you have both sides are in the same ln, you can ignore them and get -5x +3 = 2x +2 and solve for x. that is it. don't forget, the condition to "ignore " is : the same base of log. in your case, both are ln. so you can apply that trick. if they are not the same like log 5 of 6 and log 2 of 7. You can NOT